OpenStudy (anonymous):

While simplifying some math work, Peter wrote on his paper that x3 • x3 • x3 • x3 equaled x3+ 3 + 3 +3 . Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify x3 + x3 + x3 + x3?

OpenStudy (anonymous):

Is x3 supposed to be x^3 or 3x

OpenStudy (anonymous):

x^3

OpenStudy (anonymous):

x^3 * x^3 * x^3 *x^3 = x^12 Add the exponents together with multiplying exponents. x^3 + x^3 + x^3 +x^3 = 4x^3, not x^12

OpenStudy (anonymous):

So the answers are yes and then no

OpenStudy (anonymous):

thank you. would it be too much to ask for if i had you explain why ? this teacher wants complete sentences when it comes to explaining ? :/

OpenStudy (anonymous):

and i didn't even know the answer worse would be trying to explain it

OpenStudy (anonymous):

Rule of multiplying exponents vs rule of adding exponents. When multiplying two numbers with exponents with the same base (in this case x), you keep the base, and add the exponents together. (x^3 * x^2) = x^5 Rule of adding exponents with the same base (also x in this case), you keep the base and add onto it the number of terms (in this case, 4). x^2 = x*x x^3 = x*x*x (x*x)*(x*x*x) = x*x*x*x*x = x*5 (x*x) + (x*x) + (x*x) = x^2 + x^2 + x^2 = 3x^2 (You have 3 x squared)

OpenStudy (anonymous):

(x*x)*(x*x*x) = x*x*x*x*x = x^5

OpenStudy (anonymous):

thank you this really helped